# AbbIE: Autoregressive Block-Based Iterative Encoder for Efficient Sequence Modeling

### v1

@@ 1 Introduction
Through extensive experiments, we show that AbbIE consistently outperforms standard Transformers for a fixed token budget. When targeting a fixed perplexity at token amounts beyond the compute-optimal point, AbbIE approaches the FLOP cost of a standard Transformer, whereas previous recursive methods have a significantly higher FLOP cost. This makes AbbIE a computationally efficient addition to large-scale training runs, as typically modern Transformers train to well above 20×20$\times$20 × the compute-optimal token count (Allal et al., 2025b; DeepSeek-AI et al., 2024). We demonstrate that by only training on two iterations, we enable AbbIE to scale to arbitrary iteration numbers with lower perplexity and improve performance on a range of in-context learning (ICL) tasks during inference. This property allows AbbIE to match the performance of a standard Transformer, while retaining the flexibility to scale inference compute when beneficial. To our knowledge, AbbIE is the only recurrent Transformer which possesses these characteristics. They make our method readily adoptable, as it functions equivalently to a standard Transformer, unless test-time scaling is explicitly desired. In summary, the contributions of this paper are:
@@ 2.1 Design
Figure 2: Architecture of AbbIE method. The model consists of three main components: Head (green), Body (blue), and Tail (yellow). Input embeddings first pass through the Head, where they are converted from token space to concept space. The embeddings are then iterated over $r$ times $(k=0,1,...,r-1)$ by the Body. The final embeddings are then processed by the Tail and projected back into token space. $N_{h}$,$N_{b}$ and $N_{t}$ show the number of Transformers blocks in each component.
@@ 2.2 Operation
[…]
@@ 4.2 Performance scaling and computational expenditure (RQ2)
Next, we evaluate the relationship between computational cost and model performance during training. As other recurrent Transformers require many iterations to train, we find it important to demonstrate that the training costs of AbbIE are much smaller than a typical recurrent formulation and are comparable with standard Transformers. Following the methodology from Hoffmann et al. (2022), we estimate the floating point operations (FLOPs) required to reach various perplexity thresholds. Results are shown in Figure 5(a). AbbIE follows the same scaling laws as standard Transformers: lower perplexities require more computation. However, it generally incurs a higher FLOP cost. Despite this, the additional cost decreases over time. To show this we define FLOP efficiency as the ratio of AbbIE FLOPs to standard Transformer FLOPs. The gradient of FLOP efficiency in Figure 5(a) is negative. Therefore, for very long training runs, on the order of 20×20$\times$20 × to 200×200$\times$200 × the compute-optimal budget, which is common in large-scale training, we hypothesize the total computational cost of AbbIE will converge to that of the standard Transformer. This is because AbbIE maintains a consistent perplexity advantage, while the marginal improvements in perplexity become harder to achieve as training progresses. Crucially we show that AbbIE-D’s improvements in perplexity and scaling stability are valid across model and batch sizes.
@@ 4.3 Test time scaling (RQ3, RQ4)
Figure 1 shows the ICL task scaling of AbbIE-D compared to Std and 0pt(only at $r=2$, where perplexity is stable). We show that AbbIE-D’s perplexity scaling has translated into improved ICL task performance. Even when the perplexity at $r=8$ is higher than the minimum the ICL task performance continues to improve. This is important, as most recurrent Transformers’s performance plateaus at the number of training iterations. Conversely, we see that once generalization is achieved in AbbIE-D, even though perplexity begins to increase after a certain point, ICL task performance continues to improve. This also shows that the relationship between ICL and perplexity is not strict. Notably we show that AbbIE-D continues to improve performance at 4×4$\times$4 × the training iteration count, which no other general recursive transformer has shown. (RQ4)

### v2

@@ 1 Introduction
Through extensive experiments, we show that AbbIE consistently outperforms standard Transformers for a fixed token budget. When targeting a fixed perplexity at token amounts beyond the compute-optimal point, AbbIE approaches the FLOP cost of a standard Transformer, whereas previous recursive methods have a significantly higher FLOP cost. This makes AbbIE a computationally efficient addition to large-scale training runs, as typically modern Transformers train to well above $20\times$ the compute-optimal token count (Allal et al., 2025b; DeepSeek-AI et al., 2024). We demonstrate that by only training on two iterations, we enable AbbIE to scale to arbitrary iteration numbers with lower perplexity and improve performance on a range of in-context learning (ICL) tasks during inference. This property allows AbbIE to match the performance of a standard Transformer, while retaining the flexibility to scale inference compute when beneficial. To our knowledge, AbbIE is the only recurrent Transformer which possesses these characteristics. They make our method readily adoptable, as it functions equivalently to a standard Transformer, unless test-time scaling is explicitly desired. In summary, the contributions of this paper are:
@@ 2.1 Design
Figure 2: Architecture of AbbIE method. The model consists of three main components: Head (green), Body (blue), and Tail (yellow). Input embeddings first pass through the Head, where they are converted from token space to concept space. The embeddings are then iterated over $r$ times (k=0,1,,r1)(k=0,1,...,r-1) by the Body. The final embeddings are then processed by the Tail and projected back into token space. $N_{h}$$N_{b}$ and $N_{t}$ show the number of Transformers blocks in each component.
@@ 2.2 Operation
where each element in the sum depends on all the previous elements. Then, for AbbIE-C, from (3), the subsequent iteration output becomes $h_{k+1}=h_{k}+\Delta h_{k}$ while for AbbIE-D, from (4), it is $h_{k+1}=2h_{k}+\Delta h_{k}$. To achieve the same output composition, the model should learn to make $h_{k}[\texttt{AbbIE-D}]=h_{k}[\texttt{AbbIE-C}]/2$, but then this would make $\Delta h_{k}[\texttt{AbbIE-D}]$ different from $\Delta h_{k}[\texttt{AbbIE-C}]$ in an unpredictable, non-linear way.
@@ 4.2 Performance scaling and computational expenditure (RQ2)
Next, we evaluate the relationship between computational cost and model performance during training. As other recurrent Transformers require many iterations to train, we find it important to demonstrate that the training costs of AbbIE are much smaller than a typical recurrent formulation and are comparable with standard Transformers. Following the methodology from Hoffmann et al. (2022), we estimate the floating point operations (FLOPs) required to reach various perplexity thresholds. Results are shown in Figure 5(a). AbbIE follows the same scaling laws as standard Transformers: lower perplexities require more computation. However, it generally incurs a higher FLOP cost. Despite this, the additional cost decreases over time. To show this we define FLOP efficiency as the ratio of AbbIE FLOPs to standard Transformer FLOPs. The gradient of FLOP efficiency in Figure 5(a) is negative. Therefore, for very long training runs, on the order of $20\times$ to $200\times$ the compute-optimal budget, which is common in large-scale training, we hypothesize the total computational cost of AbbIE will converge to that of the standard Transformer. This is because AbbIE maintains a consistent perplexity advantage, while the marginal improvements in perplexity become harder to achieve as training progresses. Crucially we show that AbbIE-D’s improvements in perplexity and scaling stability are valid across model and batch sizes.
@@ 4.3 Test time scaling (RQ3, RQ4)
Figure 1 shows the ICL task scaling of AbbIE-D compared to Std and 0pt(only at $r=2$, where perplexity is stable). We show that AbbIE-D’s perplexity scaling has translated into improved ICL task performance. Even when the perplexity at $r=8$ is higher than the minimum the ICL task performance continues to improve. This is important, as most recurrent Transformers’s performance plateaus at the number of training iterations. Conversely, we see that once generalization is achieved in AbbIE-D, even though perplexity begins to increase after a certain point, ICL task performance continues to improve. This also shows that the relationship between ICL and perplexity is not strict. Notably we show that AbbIE-D continues to improve performance at $4\times$ the training iteration count, which no other general recursive transformer has shown. (RQ4)